OpenStudy (anonymous):

Can anyone help me with this one... Solve for x: -3|2x+6|=-12 A: x=1 and x=5 B: x=-1 and x=-5 C: x=-9 and x=3 D: No solution

jagr2713 (jagr2713):

Its b and u want explanation

OpenStudy (zzr0ck3r):

\(-3|2x+6|=-12\iff|2x+6|=4\iff2x+6=4 \text{ and } \ 2x+6=-4\iff\\ x=-1,-5 \)

jagr2713 (jagr2713):

nice zz

OpenStudy (anonymous):

Nah... Just figured out what I did wrong... But thank you anyways for offering the help... Again... :D

jagr2713 (jagr2713):

yw